Can you please help me on how to do this question, I dont understand it.
Matt bikes 50km to a friend’s home. The distance, d, in kilometres, varies directly with the time, t, in hours.
a) Find an equation relating distance and time if it takes 1.5 hours to travel 24 kilometres.
b) What does the constant of variation represent in this problem?
c)Use the equation to determine how long it will take Matt to reach his friend’s house.

Answer by stanbon(75887)

Can you please help me on how to do this question, I dont understand it.
Matt bikes 50km to a friend’s home. The distance, d, in kilometres, varies directly with the time, t, in hours.
----
k = d/t
-------------
a) Find an equation relating distance and time if it takes 1.5 hours to travel 24 kilometres.
k = = 24/1.5 = 16 km/hr
--------------------------------------------
b) What does the constant of variation represent in this problem?
Ans: rate
--------------------
c)Use the equation to determine how long it will take Matt to reach his friend’s house.
k = d/t = 16 km/hr = 50 km/t ; t = 50/16 = 3.125 hrs.
=================================
Cheers,
Stan H.